Train A & B are traveling in the same direction on parallel tracks. Train A is traveling 40mph & train B is traveling 50mph, train A passes a station at 1:20 am. If train B passes throught the same ststion at 1:32 am, at what time will train B catch up to train A?
Answer by jorel1380(3719)

1:32 a.m.-1:20 a.m.=12 minutes=1/5 hour
1/5 hr x 40 mph=8 miles
40x+8=50x
10x=8
x=8/10 hours or 48 minutes
1:32+48 mins.=2:20 a.m.
----------------------
.8 x 40=32 miles
.8 x 50=40 miles ((32+8)miles)..